# Structured state-space models are deep Wiener models

### v1

@@ Appendix C Training procedure
Like most RNN architectures, SSMs are learned via simulation error minimization, see Bianchi et al. (2017). To this end, assume that a collection of input-output sequences are collected from the plant to be identified via a suitably-designed experiment campaign. We let such a dataset be denoted as $\mathcal{D}=\big{\{}\big{(}u_{0:T}^{\{l\}},y_{0:T}^{\{l\}}\big{)}_{l\in\mathcal{I}}\big{\}}$, where the superscript {${}^{\{l\}}$ is used to index the sequences over the set =$\mathcal{I}=\{1,...,N\}$. For compactness, it is assumed that these input-output sequences have a fixed length $T$, and that the input-ouput data have been normalized (Bengio et al., 2017). The sequences are partitioned into a training set $\mathcal{I}_{\text{tr}}\subset\mathcal{I}$ and a validation set $\mathcal{I}_{\text{val}}\subset\mathcal{I}$, where $\mathcal{I}_{\text{tr}}\cup\mathcal{I}_{\text{val}}=\mathcal{I}$ and $\mathcal{I}_{\text{tr}}\cap\mathcal{I}_{\text{val}}=\emptyset$.
@@ Appendix D Hyperparameters and performances of the trained SSMs
S5R — The S5R model has $L=4$ layers with $n_{\lambda}=10$ learnable eigenvalues parametrized in continuous-time. For the initialization, the eigenvalues of the diagonal matrix $\tilde{\Lambda}_{c}$ were randomly sampled using (4.3), with $(\underaccent{\bar}{r},\bar{r})=\big{(}0.1\frac{\pi}{\tau},\frac{\pi}{\tau}\big{)}$ and $(\underaccent{\bar}{\theta},\bar{\theta})=\big{(}\frac{\pi}{6},\frac{3\pi}{4}\big{)}$. The scalar $\Gamma$ was initialized to $0.9$, whereas $\tilde{B}_{c}$ and $\tilde{C}_{c}$ were initialized with the Xavier method.

### v2

@@ Appendix C Training procedure
Like most RNN architectures, SSMs are learned via simulation error minimization, see Bianchi et al. (2017). To this end, assume that a collection of input-output sequences are collected from the plant to be identified via a suitably-designed experiment campaign. We let such a dataset be denoted as $\mathcal{D}=\big{\{}\big{(}u_{0:T}^{\{l\}},y_{0:T}^{\{l\}}\big{)}_{l\in\mathcal{I}}\big{\}}$, where the superscript {l} is used to index the sequences over the set =$\mathcal{I}=\{1,...,N\}$. For compactness, it is assumed that these input-output sequences have a fixed length $T$, and that the input-ouput data have been normalized (Bengio et al., 2017). The sequences are partitioned into a training set $\mathcal{I}_{\text{tr}}\subset\mathcal{I}$ and a validation set $\mathcal{I}_{\text{val}}\subset\mathcal{I}$, where $\mathcal{I}_{\text{tr}}\cup\mathcal{I}_{\text{val}}=\mathcal{I}$ and $\mathcal{I}_{\text{tr}}\cap\mathcal{I}_{\text{val}}=\emptyset$.
@@ Appendix D Hyperparameters and performances of the trained SSMs
S5R — The S5R model has $L=4$ layers with $n_{\lambda}=10$ learnable eigenvalues parametrized in continuous-time. For the initialization, the eigenvalues of the diagonal matrix $\tilde{\Lambda}_{c}$ were randomly sampled using (14), with $(\underaccent{\bar}{r},\bar{r})=\big{(}0.1\frac{\pi}{\tau},\frac{\pi}{\tau}\big{)}$ and $(\underaccent{\bar}{\theta},\bar{\theta})=\big{(}\frac{\pi}{6},\frac{3\pi}{4}\big{)}$. The scalar $\Gamma$ was initialized to $0.9$, whereas $\tilde{B}_{c}$ and $\tilde{C}_{c}$ were initialized with the Xavier method.